# A Model Predictive Capture Point Control Framework for Robust Humanoid Balancing via Ankle, Hip, and Stepping Strategies

### v2

@@ VI-E Simulation to Validate the Effectiveness of the Variable Weighting Method
In order to evaluate the effectiveness of the variable weighting method in the presence of disturbance, a comparative simulation with the constant weighting method was performed. The simulation setup was identical to that in Sec VI-B, where an external force was applied to the robot’s pelvis in the negative x-direction during walking in place. In this simulation, ZMP control and CAM control without stepping control were employed, and only the effects of external forces in the x-direction were analyzed. In the constant weighting method, $\mathbf{w}_{\tau}$ is set equal to $\mathbf{w}_{\tau,max}$ in the variable weighting method, aiming to restrict CAM generation for small disturbances, as in the variable weighting method.
@@
VI-F Simulation to evaluate walking performance based on different parameter selections of the stepping controller
@@
Figure 10: Planned footstep placement used in the simulation of Section VI-F (a) forward and backward walking at various speeds in the x-direction. (b) lateral walking while the robot doesn’t move in the x-direction.
@@
VI-G Simulations to Compare Robustness against Disturbances with QP-based CP controller [39]
In this section, simulations were conducted to compare the robustness against disturbances between the proposed method and state-of-the-art QP-based CP controller [39]. For the comparative simulations, the algorithm proposed by [39] was implemented in our walking control framework (see Fig. 2). In the implemented algorithm, the QP-based CP controller replaced our CP–MPC and stepping controller, while all other controllers and planners, including the ZMP controller, CAM controller, and walking pattern generator described in Section III, remained unchanged.
@@ VI-G Simulations to Compare Robustness against Disturbances with QP-based CP controller [39]
where the subscript $err$ signifies the difference between the optimization variables and the pre-designed nominal variables. The delta ZMP $\Delta\mathbf{z}=[\Delta z_{x}\,\Delta z_{y}$] denotes the difference between the desired ZMP for controlling the CP and the pre-designed reference ZMP. The cost function (VI-G) comprises the step position error term, step time error term, damping term of the CAM, and CP offset error term. The equality constraint in (VI-G) is derived from the error dynamics of the CP end-of-step dynamics in LIPFM. Detailed derivation and explanations for each equation are provided in [39].
@@
The QP formulation in (VI-G) and (VI-G) addresses the ankle strategy through delta ZMP $\Delta\mathbf{z}\in\mathbb{R}^{2}$, the hip strategy through centroidal moment $\bm{\tau}\in\mathbb{R}^{2}$, and the stepping strategy through footstep position $\mathbf{f}\in\mathbb{R}^{2}$ and step time $\gamma$. However, due to the variable coupling between $\Delta\mathbf{z}$ and $\gamma$, as well as between $\gamma$ and $\bm{\tau}$, the original equation of (VI-G) resulted in a non-linear constraint. To address this issue in [39], $\Delta\mathbf{z}$ is considered as a constant using the pre-computed value from the instantaneous CP end-of-step controller proposed in [40] without optimization. Furthermore, to avoid variable coupling between $\bm{\tau}$ and $\gamma$, a constant $\gamma^{\prime}$ was used, which corresponds to the $\gamma$ from the previous control cycle. In contrast, our approach optimizes ZMP, centroidal moment, and footstep position simultaneously using CP–MPC, and the step time is separately optimized in the stepping controller without any constant assumptions.
@@
Simulations were conducted to compare the robustness of the two algorithms against external forces. The external forces were applied to the robot during walking in place and forward walking, and the maximum impulse that the robot could withstand was analyzed using DP. The weighting parameter in (VI-G) was the same as that used in [39]. Fig. 11 shows the DP obtained when using both methods. Overall, our method exhibited more robust balancing performance, especially at 0, 30, and 330 degrees, compared to the QP-based CP controller. The overall difference in balancing performance may first be attributed to the ankle strategy. In [39], the desired ZMP for instantaneous CP control is calculated as follows,
$\mathbf{z}^{des}$ $=\mathbf{z}^{ref}+\Delta\mathbf{z}$
$=\mathbf{z}^{ref}-\frac{e^{\omega(T-t)}}{1-e^{\omega(T-t)}}\bm{\xi}_{err}.$ (26)
@@
VI-H Real Robot Experiments to Validate Robustness of the Proposed Method under External Forces
@@
Real robot experiments were conducted where the robot was subjected to external forces in both x- and y-directions during walking in place, replicating the scenarios in the simulation presented in Section VI-B. The external force was directly applied to the robot by a human using a tool, and a F/T sensor mounted on the tool was used to measure the timing and magnitude of the applied external force. The step duration includes an SSP of 0.6 s and a DSP of 0.3 s which is consistent with the simulation setup. The control parameters used in both the simulation and experiment are identical, except for the constraints of the centroidal moment (as shown in Table. I). The difference in centroidal moment constraints arises from the conservative assignment of upper body motion constraints in real robot experiments. This is done to prevent self-collisions with equipment attached to the robot when the upper body moves. Consequently, this implies that if the low-level control of the real robot is already adjusted to suit the hardware, there is no need for tuning of control parameters during sim-to-real implementation.
@@
In the first experiment, the robot was subjected to an external push of 45.95 Ns in the negative x-direction while in the right foot support. The experimental snapshots and the corresponding data are shown in Fig. 12(a). The snapshot illustrates the robot maintaining balance against external forces. As in Section VI-B, the robot successfully overcame the external force using a combination of CAM control, ZMP control, and stepping control. At approximately 6.7 s, when the external force is applied to the robot, the CP error $\xi_{err,x}$ (green line) increases in the negative x-direction, and the desired ZMP $z^{des}_{x}$ (red line) is rapidly generated in the direction that reduces the CP error in the CP–MPC. The desired ZMP is then constrained by the ZMP constraint of -9 cm, and the CP–MPC generates a footstep adjustment $\Delta f_{x,1}$ (blue line) with a peak value of approximately -14.2 cm to relax the ZMP constraint at approximately 6.9 s. Additionally, the desired centroidal moment $\tau^{des}_{y}$ (yellow line) is generated up to -7.0 Nm to further reduce the CP error at approximately 7.0 s. Despite these efforts, the robot was unable to overcome the strong disturbance. As a result, CP–MPC continuously generated footstep adjustment to reduce the CP error while the desired ZMP was generated up to the constraint until approximately 8.9 s. In conclusion, a total of three back-step controls (magenta and cyan lines) were performed, involving a movement of approximately -29.8 cm, along with step time optimization $T$ (orange line) and the robot maintained its balance. In contrast, when the QP-based CP controller proposed in [39] was used in the experiment, despite an external push of 38.7 Ns in the negative x-direction, the CP of the robot could not be recovered and the robot lost its balance.
@@
In the following experiment, the robot was subjected to an external push of 96.01 Ns in the negative y-direction during left foot support. The snapshots and the corresponding experimental data are presented in Fig. 12(b). Similar to the previous experiment, the robot maintained its balance through the three balance strategies. At approximately 7.5 s, the external force caused significant perturbation to the robot’s CP, resulting in an increase in CP error $\xi_{err,y}$ in the negative y-direction. To reduce the CP error, the desired ZMP $z^{des}_{y}$ was generated up to the ZMP constraint at approximately 7.7 s. The CP–MPC then generated a footstep adjustment $\Delta f_{y,1}$ of approximately -7.8 cm and a maximum centroidal moment $\tau_{x}^{des}$ of 7.0 Nm. Additionally, the stepping controller reduced the step time $T$ from 0.6 s to 0.43 s. Subsequently, the desired ZMP was continuously generated up to the ZMP constraint to reduce the CP error until approximately 8.5 s. However, centroidal moment and footstep adjustment were reduced after 8.0 s, and CP control was handled mainly through ZMP control. Finally, the robot maintained balance against external forces within two steps. On the other hand, when using the QP-based CP controller [39], the robot could not withstand an external force of approximately 75.43 Ns and lost balance. Overall, the magnitude of external impulses applied to the robot was similar to those in the simulation, although it was hard to make them identical due to the human factor.
@@
VI-I Real Robot Experiment to Validate Robustness of the Proposed Method on Uneven Terrain
@@
To validate the robustness of the proposed method in uneven terrain, a real robot experiment was conducted. In a scenario similar to the simulation in Section VI-C, three objects, each with a length of 15 cm and width of 20 cm, were placed in front of the robot. The distance between the first and second objects was 25 cm, and the distance between the second and third objects was 40 cm. The objects were placed alternately on the left and right sides of the robot’s center. The thickness of the objects ranged from 1.5 cm to 2.5 cm in increments of 0.5 cm. The robot walked a distance of 1.3 m with a step length of 0.1 m, and the walking duration consisted of 0.6 s of SSP and 0.3 s of DSP.
@@
Fig. 13 presents experimental snapshots and the corresponding output data. The snapshots provide a reaction of the robot to each object before and after stepping on it. While the first object was overcome by ZMP control, both the second and third objects were overcome through a combination of stepping control and CAM control in addition to ZMP control.
@@
Based on the data presented in the graph, the balancing process can be explained. During forward walking, the robot stepped onto an object with a height of 1.5 cm at the start of the 5th DSP, approximately 6.3 s. However, the robot’s CP was not significantly perturbed by the relatively low thickness of the object. As a result, the robot maintained balance through ZMP control alone until approximately 7.5 s.
@@
At the beginning of the 8th DSP, approximately 8.9 s, the robot stepped at an angle on a 2.0 cm object, leading to a significant increase in CP error in both the x- and y-directions. To reduce the CP error, CP–MPC generates the desired ZMP not only up to the ZMP constraints in both x- and y-directions but also generates footstep adjustments $\Delta f_{x,y,1}$ and the desired centroidal moment $\tau_{x,y}^{des}$. Additionally, the stepping controller adjusts the swing foot position $e^{ref}_{L,R}$ based on these footstep adjustments $\Delta f_{x,y,1}$ and the step time $T$. In contrast to the disturbance from the initial object, the three balance strategies are actively employed.
Finally, from the beginning of the 11th DSP at 11.6 s to the end of the 14th DSP at 14.6 s, the robot consistently steps on a 2.5 cm object obliquely in the y-direction. As a result, a significant disturbance occurs, particularly in the y-direction compared to the x-direction. Initially, the robot reduces CP error using only ZMP control, but notably from around 13.5 s, the three balance strategies are simultaneously employed to control substantial y-direction CP error, ultimately enabling the robot to maintain balance.
@@
Similar to the results in Section VI-C, where the robot primarily overcame disturbances in the x-direction using ZMP control with a long support foot length, in the y-direction, both stepping and CAM control were actively utilized. When using the QP-based CP controller [39], however, the robot could not withstand the disturbance from a 2.5 cm object and fell over.
VII Discussion
VII-A Hierarchical Structure of CP–MPC and Stepping Controller
VII-A1 Optimality
@@
In the proposed walking framework, the CP–MPC and the stepping controller are structured hierarchically. The CP–MPC is not iteratively re-optimized within the same control cycle based on adjusted step time $T$ and footstep position $\mathbf{f}$ from the stepping controller. In order to analyze the impact of the adjusted step time and footstep position on the optimality of CP–MPC, we re-optimized CP–MPC by incorporating the adjusted footstep $\mathbf{f}$ and step time $T$. We then analyzed the cost values (10) of CP–MPC during this re-optimization process to examine the impact of these adjustments on its optimality.
@@
However, performing more than two iterations incurs a significant computational load, making it challenging to achieve a 50 Hz operation with our 1.5 s horizon setup. Based on the simulation results, to achieve a two-time iteration, it is necessary to adjust the CP–MPC’s operation to 40 Hz and reduce the MPC horizon to 1.0 s. A shorter horizon time and a decrease in control frequency can degrade MPC performance. Consequently, this leads to weakened balancing performance against disturbances.
@@
VII-A2 Adavantages and limitations

### v3

@@ VII-E Simulation to Validate the Effectiveness of the Variable Weighting Method
In order to evaluate the effectiveness of the variable weighting method in the presence of disturbance, a comparative simulation with the constant weighting method was performed. The simulation setup was identical to that in Sec VII-B, where an external force was applied to the robot’s pelvis in the negative x-direction during walking in place. In this simulation, ZMP control and CAM control without stepping control were employed, and only the effects of external forces in the x-direction were analyzed. In the constant weighting method, $\mathbf{w}_{\tau}$ is set equal to $\mathbf{w}_{\tau,max}$ in the variable weighting method, aiming to restrict CAM generation for small disturbances, as in the variable weighting method.
@@
VII-F Simulation to evaluate walking performance based on different parameter selections of the stepping controller
@@
Figure 9: Planned footstep placement used in the simulation of Section VII-F (a) forward and backward walking at various speeds in the x-direction. (b) lateral walking while the robot doesn’t move in the x-direction.
@@
VII-G Simulations to Compare Robustness against Disturbances with QP-based CP controller [39]
In this section, simulations were conducted to compare the robustness against disturbances between the proposed method and state-of-the-art QP-based CP controller [39]. For the comparative simulations, the algorithm proposed by [39] was implemented in our walking control framework (see Fig. 1). In the implemented algorithm, the QP-based CP controller replaced our CP–MPC and stepping controller, while all other local controllers and trajectory planners described in Section III remained unchanged.
@@ VII-G Simulations to Compare Robustness against Disturbances with QP-based CP controller [39]
where the subscript $err$ signifies the difference between the optimization variables and the pre-designed nominal variables. The delta ZMP $\Delta\mathbf{z}=[\Delta z_{x}\,\Delta z_{y}$] denotes the difference between the desired ZMP for controlling the CP and the pre-designed reference ZMP. The cost function (VII-G) comprises the step position error term, step time error term, damping term of the CAM, and CP offset error term. The equality constraint in (VII-G) is derived from the error dynamics of the CP end-of-step dynamics in LIPFM. Detailed derivation and explanations for each equation are provided in [39].
@@
The QP formulation in (VII-G) and (VII-G) addresses the ankle strategy through delta ZMP $\Delta\mathbf{z}\in\mathbb{R}^{2}$, the hip strategy through centroidal moment $\bm{\tau}\in\mathbb{R}^{2}$, and the stepping strategy through footstep position $\mathbf{f}\in\mathbb{R}^{2}$ and step time $\gamma$. However, due to the variable coupling between $\Delta\mathbf{z}$ and $\gamma$, as well as between $\gamma$ and $\bm{\tau}$, the original equation of (VII-G) resulted in a non-linear constraint. To address this issue in [39], $\Delta\mathbf{z}$ is considered as a constant using the pre-computed value from the instantaneous CP end-of-step controller proposed in [40] without optimization. Furthermore, to avoid variable coupling between $\bm{\tau}$ and $\gamma$, a constant $\gamma^{\prime}$ was used, which corresponds to the $\gamma$ from the previous control cycle. In contrast, our approach optimizes ZMP, centroidal moment, and footstep position simultaneously using CP–MPC, and the step time is separately optimized in the stepping controller without any constant assumptions.
@@
Simulations were conducted to compare the robustness of the two algorithms against external forces. The external forces were applied to the robot during walking in place and forward walking, and the maximum impulse that the robot could withstand was analyzed using DP. The weighting parameter in (VII-G) was the same as that used in [39]. Fig. 10 shows the DP obtained when using both methods. Overall, our method exhibited more robust balancing performance, especially at 0, 30, and 330 degrees, compared to the QP-based CP controller. The overall difference in balancing performance may first be attributed to the ankle strategy. In [39], the desired ZMP for instantaneous CP control is calculated as follows,
$\mathbf{z}^{des}$ $=\mathbf{z}^{ref}+\Delta\mathbf{z}=\mathbf{z}^{ref}-\frac{e^{\omega(T-t)}}{1-e^{\omega(T-t)}}\bm{\xi}_{err}.$ (27)
@@
VII-H Real Robot Experiments to Validate Robustness of the Proposed Method under External Forces
@@
Real robot experiments were conducted where the robot was subjected to external forces in both x- and y-directions, replicating the scenarios in the simulation presented in Section VII-B. The external force was applied directly to the robot by a human using a tool equipped with an F/T sensor to measure the timing and magnitude of the force. The step duration includes an SSP of 0.6 s and a DSP of 0.3 s which is consistent with the simulation setup. The control parameters used in both the simulation and experiment are identical, except for the constraints of the centroidal moment (as shown in Table. I). The difference in centroidal moment constraints results from conservative upper body motion limits in real robot experiments, applied to prevent collisions with equipment attached to the robot. Consequently, this implies that if the low-level control of the real robot is already adjusted to suit the hardware, there is no need for tuning of control parameters during sim-to-real implementation.
@@
In the first experiment, the robot was subjected to an external push of 44.98 Ns in the negative x-direction while in the right foot support. The experimental snapshots and the corresponding data are shown in Fig. 11(a). The snapshot illustrates the robot maintaining balance against external forces. As in Section VII-B, the robot successfully overcame the external force using a combination of ZMP control, CAM control, and stepping control. At approximately 6.5 s, when the external force is applied to the robot, the CP error $\xi_{err,x}$ (green line) increases in the negative x-direction, and the desired ZMP $z^{des}_{x}$ (red line) is rapidly generated in the direction that reduces the CP error in the CP–MPC. The desired ZMP is then constrained by the ZMP constraint of -9 cm, and the CP–MPC generates a footstep adjustment $\Delta f_{x,1}$ (blue line) with a peak value of approximately -11.5 cm to relax the ZMP constraint at approximately 7.0 s. Additionally, the desired centroidal moment $\tau^{des}_{y}$ (yellow line) is generated up to -7.0 Nm to further reduce the CP error at approximately 7.1 s. Despite these efforts, the robot was unable to overcome the strong disturbance. As a result, CP–MPC continuously generated footstep adjustment to reduce the CP error while the desired ZMP was generated up to the constraint until approximately 8.5 s. In conclusion, a total of two back-step controls (magenta and cyan lines) were performed, involving a movement of approximately -15 cm, along with step time optimization $T$ (orange line) and the robot maintained its balance. In contrast, when the QP-based CP controller proposed in [39] was used in the experiment, despite an external push of 38.7 Ns in the negative x-direction, the CP of the robot could not be recovered and the robot lost its balance.
@@
In the following experiment, the robot was subjected to an external push of 94.71 Ns in the negative y-direction during left foot support. The snapshots and the corresponding experimental data are presented in Fig. 11(b). At approximately 7.4 s, the external force caused significant perturbation to the robot’s CP, resulting in an increase in CP error $\xi_{err,y}$ in the negative y-direction. To reduce the CP error, the desired ZMP $z^{des}_{y}$ was generated up to the ZMP constraint at approximately 7.6 s. The CP–MPC then generated a footstep adjustment $\Delta f_{y,1}$ of approximately -8.6 cm and a maximum centroidal moment $\tau_{x}^{des}$ of 7.0 Nm. Additionally, the stepping controller reduced the step time $T$ from 0.6 s to 0.48 s. Subsequently, the desired ZMP was continuously generated up to the ZMP constraint to reduce the CP error until approximately 8.1 s. However, centroidal moment and footstep adjustment were reduced after 7.8 s, and CP control was handled mainly through ZMP control. Finally, the robot maintained balance against external forces. On the other hand, when using the QP-based CP controller [39], the robot could not withstand an external force of approximately 75.43 Ns and lost balance. Overall, the magnitude of external impulses applied to the robot was similar to those in the simulation, although it was hard to make them identical due to the human factor.
@@
VII-I Real Robot Experiment to Validate Robustness of the Proposed Method on Uneven Terrain
@@
To validate the robustness of the proposed method in uneven terrain, a real robot experiment was conducted. In a scenario similar to the simulation in Section VII-C, three objects, each with a length of 15 cm and width of 20 cm, were placed in front of the robot. The distance between the first and second objects was 25 cm, and the distance between the second and third objects was 40 cm. The objects were placed alternately on the left and right sides of the robot’s center. The thickness of the objects ranged from 2.0 cm to 3.0 cm in increments of 0.5 cm. The robot walked a distance of 1.3 m with a step length of 0.1 m, and the walking duration consisted of 0.6 s of SSP and 0.3 s of DSP.
@@
Fig. 12 presents experimental snapshots and the corresponding output data. The snapshots provide a reaction of the robot to each object before and after stepping on it. The first object was primarily overcome using ZMP control, while the second and third objects were addressed through a combination of stepping and CAM control in addition to ZMP control.
@@
Based on the data presented in the graph, the balancing process can be explained. During forward walking, the robot stepped onto a 2.0 cm object at the start of the 5th DSP, approximately 6.3 s. However, the robot’s CP was not significantly perturbed, allowing it to overcome the obstacle primarily through ZMP control until approximately 7.5 s.
@@
At the start of the 8th DSP, approximately 9.0 s, the robot stepped at an angle on a 2.5 cm object, resulting in a significant increase in CP error in both the x- and y-directions until the 11th SSP. To reduce this error, CP–MPC not only generates the ZMP up to the ZMP constraints in both x- and y-directions but also generates footstep adjustments $\Delta f_{x,y,1}$ and the desired centroidal moment $\tau_{x,y}^{des}$. The stepping controller then adjusts the swing foot position $e^{ref}_{L,R}$ based on these footstep adjustments $\Delta f_{x,y,1}$ and the step time $T$. Unlike the disturbance from the initial object, all three balance strategies are actively employed.
Immediately after the balance control performed until the 11th SSP, from the start of the 11th DSP at 11.7 s to the end of the 14th DSP at 14.6 s, the robot continuously stepped obliquely on a 3.0 cm object in the y-direction. This caused a significantly larger disturbance in the y-direction compared to the x-direction. In response, the robot actively employed the three balance strategies to control the y-direction CP error, ultimately enabling it to maintain balance.
@@
Similar to the results in Section VII-C, where the robot primarily overcame disturbances in the x-direction using ZMP control with a long support foot length, in the y-direction, both stepping and CAM control were actively utilized. When using the QP-based CP controller [39], however, the robot could not withstand the disturbance from a 3.0 cm object and fell over.
VIII Discussion
VIII-A Hierarchical Structure of CP–MPC and Stepping Controller
VIII-A1 Optimality
@@
In the proposed walking framework, the CP–MPC and the stepping controller are structured hierarchically. The CP–MPC is not re-optimized within the same control cycle based on adjusted step time $T$ and footstep position $\mathbf{f}$ from the stepping controller. In order to analyze the impact of the adjusted step time and footstep position on the optimality of CP–MPC, we re-optimized CP–MPC by incorporating the adjusted footstep $\mathbf{f}$ and step time $T$. We then analyzed the cost values (10) of CP–MPC during this re-optimization process to examine the impact of these adjustments on its optimality.
@@
However, more than two iterations incurs a significant computational load, making it difficult to maintain a 50 Hz operation with our 1.5 s horizon setup. To achieve a two-time iteration, it is necessary to adjust the CP–MPC’s operation to 40 Hz and reduce the MPC horizon to 1.0 s.
@@
Consequently, the shortened horizon and reduced control frequency negatively impacted the robot’s robustness during walking. From this perspective, as suggested in [53], implementing an intrinsically stable MPC that is less dependent on horizon length could be one of the promising ideas for improving the overall performance of MPC.
VIII-A2 Adavantages and limitations